### v1

Abstract
We present an implementation and a brief experimental analysis of the deterministic algorithm proposed by Duan et al. (2025) for the Single-Source Shortest Path (SSSP) problem, which achieves the best known asymptotic upper bound in the comparison-addition model, with running time $O(m\log^{2/3}n)$. We provide a faithful C++ implementation of this algorithm, following all structural details described in the original paper, and compare its empirical performance with the classical Dijkstra’s algorithm using binary heaps. The experiments were conducted on both synthetic sparse random graphs and real-world road network instances from the DIMACS benchmark. Our results show that, despite its superior asymptotic complexity, the new algorithm presents significantly larger constant factors, making Dijkstra’s algorithm faster for all tested sparse graph sizes, including instances with tens of millions of vertices. Our implementation achieves $O(m\log^{2/3}n)$ expected time, due to the use of hash tables, and some possibilities for making it worst-case are being considered. (This is a ongoing work.)
keywords:
Single Source Shortest Path, Dijkstra’s Algorithm, Graph, Algorithms, Experimental Analysis, Time Complexity
category:
1 Introduction
Let $G=(V,E)$, $|V|=n,|E|=m$, be a connected graph with a weight function $w:E\rightarrow\mathbb{R}_{\geq 0}$ and a source vertex $s\in V$. The Single Source Shortest Path Problem (SSSP) consists of determining, for each vertex $v\in V$, the minimum distance $d(v)=d(s,v)$ of a path in $G$ that starts at $s$ and ends at $v$.
As one of the most classical algorithmic problems in graph theory, the SSSP is widely studied in the literature. Until the end of the last century, the algorithm proposed by dljkstra1959note, was still considered the state of the art for solving the SSSP. When combined with an efficient priority queue, like binary heap (williams_algorithm_1964), Dijkstra’s algorithm solves the problem with a time complexity of $O((n+m)\log n)$. In haeupler2024universal, haeupler2024universal showed that Dijkstra’s algorithm is universally optimal for the natural problem of ordering the vertices according to their distances $d(v)$ from the source $s$. In sparse graphs, where $m=O(n)$, this complexity matches the sorting barrier in the comparison-addition computation model.
To overcome the bottleneck associated with vertex ordering, some works have proposed algorithms with lower time complexity by avoiding sorting the vertices. In the RAM computation model, thorup1999undirected presented an algorithm with time complexity $O(m)$ for undirected graphs whose weights fit into a machine word, and later (thorup2004integer) extended the approach to directed graphs, achieving $O(m+n\log\log n)$. In the comparison-addition model, duan2023randomized proposed a randomized algorithm for undirected graphs with time complexity $O(m\sqrt{\log n\log\log n})$.
Recently, duan2025breaking presented the first deterministic algorithm with $o(n\log n)$ complexity for sparse graphs, achieving an upper bound of $O(m\log^{2/3}n)$.
This work aims to empirically evaluate the deterministic algorithm proposed by duan2025breaking and compare its practical performance with Dijkstra’s classical algorithm. We implement both algorithms in C++ (Dijkstra’s algorithm using a binary heap, and Duan’s algorithm faithfully following the implementation details described in their paper).
We acknowledge the recent Rust implementation by valko2025outperforming and their evaluation on Lightning Network topologies. Their work provides a practical approach, using sorting as their selection algorithm. This differs from the theoretical proposal in duan2025breaking, which assumes a worst-case linear-time selection algorithm. As a result of this implementation choice, the algorithm’s complexity aligns with that of Dijkstra, though with higher constant factors. This implementation was evaluated on graphs of up to approximately 15k nodes, offering useful insights for networks of that particular size and topology.
To the best of our knowledge, this is the first work to present a faithful implementation of Duan’s algorithm and provide a practical performance analysis on random and road networks up to approximately 10M vertices.
The remainder of this paper is organized as follows: Section 2 outlines the main ideas of the new algorithm; Section 3 analyzes the expected behavior of the comparison; Section 4 presents key implementation details; Section 5 describes the experimental setup; Section 6 discusses the obtained results; and Section 7 concludes the paper.
2 Duan et al. (2025)’s algorithm
The algorithm proposed by duan2025breaking addresses an extension of the SSSP called Bounded Multi-source Shortest Path (BMSSP). In BMSSP, the goal is to compute, for each vertex $v\in V$, the minimum distance $d(v)$ from some vertex $u\in S\subseteq V$, subject to the constraint $d(v)<B$. When $S=\{s\}$ and $B=\infty$, BMSSP reduces to the standard SSSP.
The algorithm uses two parameters, $k$ and $t$, and is based on a divide-and-conquer approach over the vertex set. The vertex set is recursively divided into $2^{t}$ roughly equal parts across $O((\log n)/t)$ recursion levels, until reaching the base case where a subproblem contains a single vertex $x$. At the base case, Dijkstra’s algorithm is executed to compute the shortest distances from $x$ to its $k$ nearest vertices.
The algorithm faces two main bottlenecks:
  1. 1.
    selecting the new set $S$ to connect sequential subproblems in a partition, and
  2. 2.
    controlling the size of $S$ to prevent degeneration of the overall complexity.
To address the first bottleneck, duan2025breaking use a specialized data structure to select the $2^{(l-1)t}$ vertices with the smallest distance in each partition at recursion level $l$. This data structure has been fully implemented in this work. To address the second bottleneck, the set $S$ is reduced to a smaller set of pivots $P\subseteq S$ that suffice as sources. This reduction is performed through $k$ iterations of a Bellman-Ford-like algorithm (bellman1958routing).
The parameters $k$ and $t$ are chosen to balance the cost of the Bellman-Ford iterations with the recursion depth, resulting in a final time complexity of $O(m\log^{2/3}n)$.
3 Constant Analysis
Asymptotic complexity characterizes an algorithm’s growth rate by focusing on the dominant term of its running time and ignoring constants and lower-order terms. However, in practice, these constants can significantly impact performance on real computers. In this section, we present an analysis of the constants in the BMSSP algorithm to allow a more precise comparison with Dijkstra’s algorithm. (For simplicity, we focus only on the dominant term and neglect the other ones.)
For sparse graphs, Dijkstra’s algorithm has time complexity $O(n\log n)$, which can be approximated as $c_{1}\cdot n\log n$, where $c_{1}$ is a constant reflecting the implementation overhead.
Similarly, the BMSSP algorithm has complexity $O(n\log^{2/3}n)$, which we approximate as $c_{2}\cdot n\log^{2/3}n$, where $c_{2}$ is a constant reflecting the implementation overhead.
Since $O(n\log n)$ grows faster than $O(n\log^{2/3}n)$, there exists a threshold $n_{0}$ beyond which BMSSP will outperform Dijkstra. Formally, we define
$n_{0}=\min\left\{{n\mid c_{2}\,n\,\log^{2/3}n<c_{1}\,n\,\log n}\right\},$
which simplifies to
$n_{0}=\min\left\{{n\mid\log^{1/3}n>c_{2}/c_{1}}\right\}.$
Assuming different ratios $c_{2}/c_{1}$ (that is, assuming different rates at which the BMSSP constant is greater than Dijkstra’s) the threshold values $n_{0}$ are as follows:
  • $c_{2}/c_{1}=2$: BMSSP is faster for $n>64$
  • $c_{2}/c_{1}=3$: BMSSP is faster for $n\gtrsim 10^{8}$
  • $c_{2}/c_{1}=4$: BMSSP is faster for $n\gtrsim 10^{19}$
  • $c_{2}/c_{1}=5$: BMSSP is faster for $n\gtrsim 10^{38}$
  • $c_{2}/c_{1}=6$: BMSSP is faster for $n\gtrsim 10^{65}$
  • $c_{2}/c_{1}=7$: BMSSP is faster for $n\gtrsim 10^{103}$
This analysis predicts that even if the constant in a BMSSP implementation is only five times larger than that of Dijkstra’s algorithm, Dijkstra will remain faster for almost all practical graph sizes.
4 Implementation Details
We provide implementations for both algorithms discussed in this work. All implementations are written in C++20, compiled with g++ using the -O3 optimization flag. The code is publicly available at github.com/lcs147/bmssp.
Dijkstra’s algorithm was implemented using the binary heap provided by the C++ standard library (priority_queue).
The algorithm of duan2025breaking was implemented faithfully according to the description in their paper. The original algorithm is designed to run in $O(m\log^{2/3}n)$ worst-case time. Our implementation achieves $O(m\log^{2/3}n)$ expected time, due to the use of hash tables for certain set operations (such as union, membership tests, and duplicate removal). The paper does not specify the data structures for these operations, and we found no alternative that would preserve both the $O(m)$ memory usage and the $O(m\log^{2/3}n)$ time complexity. Therefore, hash tables were chosen as a practical solution for implementing these operations efficiently.
5 Experiments
All experiments were conducted on a computer with 32 GB of memory and an Intel Core i5-10400F processor running at 2.90 GHz, under Linux Mint 21.3 Cinnamon.
We tested two types of instances:
  • Sparse Random Graphs: randomly generated graphs with sizes $2^{7},2^{8},\dots,2^{24}$. These graphs have a mean out-degree of 3 and a maximum out-degree of 4. The vertices were numbered from 1 to $n$, and it is guaranteed that vertex 1 can reach all other vertices. The code used to generate these instances is available in the repository.
  • Road Graphs: 12 graphs representing road networks in different regions of the United States, obtained from the 9th DIMACS Implementation Challenge (dimacs). Edge weights correspond to the average time required to travel along the respective street. Since these are road networks, all graphs are sparse. The number of vertices and edges for each instance is reported in Table 1.
For all instances, shortest paths were computed using vertex 1 as the source. Each algorithm was executed five times independently on each instance to reduce the influence of external factors such as operating system variations or timing fluctuations. The average execution time, measured in milliseconds, was used as the main performance metric. The results are discussed in Section 6.
6 Results
6.1 Sparse Random Graphs
Figure 1 shows the average execution time (on logarithmic scale) for the randomly generated graphs. Graphs with up to $2^{24}$ vertices were tested, which was the maximum size the BMSSP algorithm could handle due to the machine memory limitations.
Figure 1: Execution time (average over multiple runs) for randomly generated sparse graphs of increasing size, plotted on a log-log scale.
Figure 2 illustrates the ratio of execution times between the BMSSP and Dijkstra algorithms on random graphs. For the smallest graph size (128 vertices), BMSSP is substantially slower, measuring approximately 22 times slower than Dijkstra. As the graph size increases, however, this performance disparity significantly narrows, with the ratio dropping to approximately seven. This decreasing ratio indicates that BMSSP becomes relatively more efficient as the input size grows. Overall, across all random graph instances, BMSSP maintains an average execution time approximately nine times slower than Dijkstra.
To check if the implementation matches the expected behavior of the algorithm, we can use the asymptotic complexities of the algorithm to simulate the same ratio as the graph size increases:
$f(n)=\frac{\text{Time(BMSSP)}}{\text{Time(Dijkstra)}}=\frac{n\log^{2/3}n}{n\log n}=\log^{-1/3}n.$
Figure 3 illustrates the expected ratio based on asymptotic analysis, and it is consistent with the ratio obtained experimentally.
Figure 2: Experimental ratio of BMSSP to Dijkstra execution times on random sparse graphs.
Figure 3: Theoretical ratio of BMSSP to Dijkstra execution times based on asymptotic analysis.
6.2 Road Graphs
Table 1 reports execution times for the 12 road network instances, and Figure 4 illustrates the growth of execution time with increasing graph size. Dijkstra’s algorithm was on average approximately seven times faster than the BMSSP algorithm.
While this ratio (seven) is lower than the ratio observed in the random graph experiments (nine), this result is expected. The road network instances start significantly larger, with the smallest network already exceeding 200,000 vertices (compared to 128 vertices in the smallest random graph). This larger scale provides an environment where the overhead of the BMSSP algorithm is relatively less impactful, thus reducing the performance gap between it and Dijkstra’s algorithm.
Table 1: Execution times (in milliseconds) of our implementations of Dijkstra and Duan et al. algorithms on 12 USA road network instances.
Instance $n$ $m$ Time (ms) Ratio
dljkstra1959note duan2025breaking
New York City 264,346 733,846 32.490 250.030 7.696
San Francisco Bay Area 321,270 800,172 38.803 297.297 7.662
Colorado 435,666 1,057,066 52.174 416.865 7.990
Florida 1,070,376 2,712,798 127.588 958.971 7.516
Northwest USA 1,207,945 2,840,208 155.193 1067.577 6.879
Northeast USA 1,524,453 3,897,636 228.756 1450.339 6.340
California and Nevada 1,890,815 4,657,742 257.454 1719.994 6.681
Great Lakes 2,758,119 6,885,658 376.873 2674.383 7.096
Eastern USA 3,598,623 8,778,114 553.878 3718.830 6.714
Western USA 6,262,104 15,248,146 1013.046 6903.403 6.815
Central USA 14,081,816 34,292,496 3346.156 18195.013 5.438
Full USA 23,947,347 58,333,344 4562.473 29345.319 6.432
Figure 4: Execution time versus graph size for the USA road network instances, plotted on log-log scale.
6.3 General analysis
After running the tests, it was observed that the classical algorithm of dljkstra1959note achieved superior performance in all evaluated instances. Although, in theory, the algorithm proposed by duan2025breaking has a better asymptotic upper bound, its implementation is considerably more complex, involving multiple recursive calls and nuances that introduce a large constant factor in the running time. In practice, this overhead causes the algorithm to perform worse than Dijkstra’s algorithm, which is simple, direct, and highly efficient—especially when using, for example, a binary heap structure implemented directly in an array.
If we assume that the constant of our BMSSP implementation is seven times higher than Dijkstra’s ($c_{2}/c_{1}=7$), a ratio derived from the road network experiment, then the same analysis used in Section 3 predicts that our implementation will only be faster when the graph size surpasses $10^{103}$.
These results indicate that the theoretical improvement obtained through asymptotic analysis does not necessarily translate into practical gains, due to the large hidden constants and structural complexity of the new algorithm. In most real-world environments, using Dijkstra’s algorithm is therefore substantially more efficient.
7 Concluding Remarks
We presented an experimental analysis comparing Dijkstra’s classical algorithm (dljkstra1959note) with the recent algorithm by duan2025breaking, which establishes the best known deterministic asymptotic SSSP upper bound in the comparison-addition model.
Our experiments show that Dijkstra’s algorithm remains the preferred choice for graphs with up to ten million vertices, covering most practical applications. Also, we predict that Dijkstra’s should be the preferred choice for graphs much bigger than that. (As discussed in Section 3, its large hidden constants limit its practical efficiency, despite its superior asymptotic bound.)
Dijkstra’s algorithm remains the standard for real-world graph problems. However, the work of duan2025breaking illustrates that improvements may be achievable, inspiring the exploration of algorithms that combine strong asymptotic performance with low practical overhead.