### v2

Abstract
We present an implementation and a brief experimental analysis of the deterministic algorithm proposed by Duan et al. (2025) for the Single-Source Shortest Path (SSSP) problem, which achieves the best known asymptotic upper bound in the comparison-addition model, with running time $O(m\log^{2/3}n)$. We provide a faithful C++ implementation of this algorithm, following all structural details described in the original paper, and compare its empirical performance with the classical Dijkstra’s algorithm using binary heaps. The experiments were conducted on both synthetic sparse random graphs and real-world road network instances from the DIMACS benchmark. Our implementation achieves the proposed running time in the worst case. However, our results show that, despite the new algorithm’s superior asymptotic complexity, its large constant factors significantly affect its practical performance, making Dijkstra’s algorithm faster for all tested sparse graph sizes, including instances with tens of millions of vertices. (This is a work in progress.)
keywords:
Algorithms, Dijkstra’s Algorithm, Experimental Analysis, Heap and Hash Data Structures, Single Source Shortest Path, Time Complexity.
category:
1 Introduction
Let $G=(V,E)$, $|V|=n,|E|=m$, be a connected graph with a weight function $w:E\rightarrow\mathbb{R}_{\geq 0}$ and a source vertex $s\in V$. The Single Source Shortest Path Problem (SSSP) consists of determining, for each vertex $v\in V$, the minimum distance $d(v)=d(s,v)$ of a path in $G$ that starts at $s$ and ends at $v$.
As one of the most classical algorithmic problems in graph theory, the SSSP is widely studied in the literature. Until the end of the last century, the algorithm proposed by dljkstra1959note, was still considered the state of the art for solving the SSSP. When combined with an efficient priority queue, like binary heap (williams_algorithm_1964), Dijkstra’s algorithm solves the problem with a time complexity of $O((n+m)\log n)$. In haeupler2024universal, haeupler2024universal showed that Dijkstra’s algorithm is universally optimal for the natural problem of ordering the vertices according to their distances $d(v)$ from the source $s$. In sparse graphs, where $m=O(n)$, this complexity matches the sorting barrier in the comparison-addition computation model.
To overcome the bottleneck associated with vertex ordering, some works have proposed algorithms with lower time complexity by avoiding sorting the vertices. In the RAM computation model, thorup1999undirected presented an algorithm with time complexity $O(m)$ for undirected graphs whose weights fit into a machine word, and later (thorup2004integer) extended the approach to directed graphs, achieving $O(m+n\log\log n)$. In the comparison-addition model, duan2023randomized proposed a randomized algorithm for undirected graphs with time complexity $O(m\sqrt{\log n\log\log n})$.
Recently, duan2025breaking presented the first deterministic algorithm with $o(n\log n)$ complexity for sparse graphs, achieving an upper bound of $O(m\log^{2/3}n)$.
This work aims to empirically evaluate the deterministic algorithm proposed by duan2025breaking and compare its practical performance with Dijkstra’s classical algorithm. We implement both algorithms in C++ (Dijkstra’s algorithm using a binary heap, and duan2025breaking algorithm faithfully following the implementation details described in their paper).
We acknowledge the recent Rust implementation by valko2025outperforming and their evaluation on Lightning Network topologies. Their work provides a practical approach, using sorting as their selection algorithm. This differs from the theoretical proposal in duan2025breaking, which assumes a worst-case linear-time selection algorithm. As a result of this implementation choice, the algorithm’s complexity aligns with that of Dijkstra, though with higher constant factors. This implementation was evaluated on graphs of up to approximately 15k nodes, offering useful insights for networks of that particular size and topology.
During the development of this work, we also became aware that makowski2025bmsspy are developing a Python package that aims to faithfully implement the algorithm proposed by duan2025breaking. They did experiments on grid and geospatial networks to compare their implementations with other Python packages which implement Dijkstra’s algorithm.
To the best of our knowledge, this is the first work to present a faithful implementation of the duan2025breaking algorithm and to provide a practical performance analysis on random and road networks up to approximately 10 million vertices.
The remainder of this paper is organized as follows: Section 2 outlines the main ideas of the new algorithm; Section 3 analyzes the expected behavior of the comparison; Section 4 presents key implementation details; Section 5 describes the experimental setup; Section 6 discusses the obtained results; and Section 7 concludes the paper.
2 Duan et al. (2025)’s Algorithm
The algorithm proposed by duan2025breaking addresses an extension of the SSSP called Bounded Multi-source Shortest Path (BMSSP). In BMSSP, the goal is to compute, for each vertex $v\in V$, the minimum distance $d(v)$ from some vertex $u\in S\subseteq V$, subject to the constraint $d(v)<B$. When $S=\{s\}$ and $B=\infty$, BMSSP reduces to the standard SSSP.
The algorithm uses two parameters, $k$ and $t$, and is based on a divide-and-conquer approach over the vertex set. The vertex set is recursively divided into $2^{t}$ roughly equal parts across $O((\log n)/t)$ recursion levels, until reaching the base case where a subproblem contains a single vertex $x$. In the base case, Dijkstra’s algorithm is executed to compute the shortest distances from $x$ to its $k$ nearest vertices.
The algorithm faces two main bottlenecks:
  1. 1.
    selecting the new set $S$ to connect sequential subproblems in a partition, and
  2. 2.
    controlling the size of $S$ to prevent degeneration of the overall complexity.
To address the first bottleneck, duan2025breaking use a specialized data structure to select the $2^{(l-1)t}$ vertices with the smallest distance in each partition at recursion level $l$. This data structure has been fully implemented in this work. To address the second bottleneck, the set $S$ is reduced to a smaller set of pivots $P\subseteq S$ that suffice as sources. This reduction is performed through $k$ iterations of a Bellman-Ford-like algorithm (bellman1958routing).
The parameters $k$ and $t$ are chosen to balance the cost of the Bellman-Ford iterations with the recursion depth, resulting in a final time complexity of $O(m\log^{2/3}n)$.
3 Asymptotic versus Empirical Analysis under Big Constants
Asymptotic complexity characterizes an algorithm’s growth rate by focusing on the dominant term of its running time and ignoring constants and lower-order terms. However, in practice, these constants can significantly impact performance on real computers. In this section, we present an analysis of the constants in the BMSSP algorithm to allow a more precise comparison with Dijkstra’s algorithm. (For simplicity, we focus only on the dominant term and neglect the other ones.)
For sparse graphs, Dijkstra’s algorithm has time complexity $O(n\log n)$, which can be approximated as $c_{1}\cdot n\log n$, where $c_{1}$ is a constant reflecting the implementation overhead.
Similarly, the BMSSP algorithm has complexity $O(n\log^{2/3}n)$, which we approximate as $c_{2}\cdot n\log^{2/3}n$, where $c_{2}$ is a constant reflecting the implementation overhead.
Since $O(n\log n)$ grows faster than $O(n\log^{2/3}n)$, there exists a threshold $n_{0}$ beyond which BMSSP will outperform Dijkstra. Formally, we define
$n_{0}=\min\left\{{n\mid c_{2}\,n\,\log^{2/3}n<c_{1}\,n\,\log n}\right\},$
which simplifies to
$n_{0}=\min\left\{{n\mid\log^{1/3}n>c_{2}/c_{1}}\right\}.$
Assuming different ratios $c_{2}/c_{1}$ (that is, assuming different rates at which the BMSSP constant is greater than Dijkstra’s) the threshold values $n_{0}$ are as follows:
  • $c_{2}/c_{1}=2$: BMSSP is faster for $n>64$
  • $c_{2}/c_{1}=3$: BMSSP is faster for $n\gtrsim 10^{8}$
  • $c_{2}/c_{1}=4$: BMSSP is faster for $n\gtrsim 10^{19}$
  • $c_{2}/c_{1}=5$: BMSSP is faster for $n\gtrsim 10^{38}$
  • $c_{2}/c_{1}=6$: BMSSP is faster for $n\gtrsim 10^{65}$
  • $c_{2}/c_{1}=7$: BMSSP is faster for $n\gtrsim 10^{103}$
This analysis predicts that even if the constant in a BMSSP implementation is only five times larger than that of Dijkstra’s algorithm, Dijkstra will remain faster for almost all practical graph sizes.
4 Implementation Details
We provide implementations for both algorithms discussed in this work. All implementations are written in C++20, compiled with g++ using the -O3 optimization flag. The code is publicly available at https://github.com/lcs147/bmssp.
The algorithm proposed by duan2025breaking requires that all simple paths have distinct lengths. To satisfy this requirement in a general setting, we must break ties consistently when paths have equal length. As described in the original paper, we represent the distance to a vertex $v$ as the tuple
$\hat{d}[v]=\big(\text{length}[v],\ \text{hops}[v],\ v,\ \text{pred}[v]\big).$
Here, $\text{length}[v]$ is the total weight of the current shortest path to $v$, $\text{hops}[v]$ is the number of edges in this path, and $\text{pred}[v]$ is the predecessor of $v$ in this path. Paths are compared using the lexicographic order on these tuples, which guarantees a unique ordering even if their length values are identical.
Dijkstra’s algorithm was implemented using the binary heap provided by the C++ standard library (std::priority_queue). We selected this data structure because binary heaps are often more efficient in practice than Fibonacci heaps (cherkassky1996shortest).
Our implementation closely follows the algorithm proposed by duan2025breaking. We implemented all three subroutines and the core data structure as specified, ensuring that the theoretical worst-case time complexity of $O(m\log^{2/3}n)$ is preserved. Below, we list some implementation details:
Additional memory usage
The block-based data structure in the original paper stores at most $N$ key-value pairs. It must support finding any key and its associated data in $O(1)$ time.
Instead of using a hash table to find the keys (which provides expected $O(1)$ operations), we employ global direct-address tables (DATs) of size $O(n)$, where $n$ is the number of vertices. Naively, each recursive call of BMSSP (Algorithm 3) would require its own DAT, leading to an infeasible number of tables. However, due to the sequential nature of the divide-and-conquer recursion, we can reuse DATs across calls. Specifically, we maintain only one DAT per recursion level, as recursive calls at the same level are processed sequentially and can share the same memory. With $O(\log^{1/3}n)$ recursion levels and a DAT of size $\Theta(n)$ per level, the total space complexity becomes $\Theta(n\log^{1/3}n)$.
Set operations
To guarantee $O(1)$ worst-case operations for set operations (such as union, membership tests, and duplicate removal), we consistently use global direct-address tables (DATs) throughout our implementation. However, unlike in the bucket data structure, the other DATs do not increase the asymptotic space complexity.
FindPivots forest building.
In the FindPivots subroutine, we do not explicitly build the directed forest $F$. Since we only need to identify roots whose trees contain at least $k$ vertices, we maintain an array $\text{root}[v]$ that stores the root of the tree to which vertex $v$ belongs. Whenever an edge $(u,v)$ is relaxed, we update $\text{root}[v]:=\text{root}[u]$. This approach allows us to efficiently determine the final set of pivots without the overhead of building the full forest and traversing any vertex.
Disjointness of $U_{i}$ sets
Experimentally, we encountered a subtle discrepancy with Remark 3.8 in duan2025breaking, which asserts that the sets $U_{i}$ in a recursive call of BMSSP (Algorithm 3) are disjoint. This property may fail. Consider a vertex $x$ that enters the data structure $D$ with an initial distance $d_{\text{old}}$. If $x$ is later marked as complete by a recursive call with a shorter distance $d_{\text{new}}$ (so that $x\in U_{i}$) but is never extracted from $D$, it remains in $D$ with the outdated key $d_{\text{old}}$. When $x$ is eventually extracted (still with key $d_{\text{old}}$), it becomes a frontier vertex, causing its shortest‑path subtree to be rediscovered with the correct distance $d_{\text{new}}$. Consequently, $x$ and its subtree may be assigned to another set $U_{j}$, violating disjointness.
We restore the disjointness property by removing a vertex from $D$ as soon as it is marked complete. Specifically, before line 15 of subroutine BMSSP (Algorithm 3), we perform $D.\mathtt{erase}(u)$. Assuming $\mathtt{erase}$ has complexity $O(\max\{1,\log(N/M)\})$ when the element is present and $O(1)$ otherwise, the asymptotic complexity of the algorithm is preserved. With this modification, the implementation satisfies the disjointness claim of Remark 3.8.
5 Experiments
All experiments were conducted on a computer with 32 GB of memory and an Intel Core i5-10400F processor running at 2.90 GHz, under Linux Mint 21.3 Cinnamon.
We tested two types of instances:
  • Sparse Random Graphs: randomly generated graphs with sizes $2^{7},2^{8},\dots,2^{25}$. These graphs have a mean out-degree of 3 and a maximum out-degree of 4. The vertices were numbered from 1 to $n$, and it is guaranteed that vertex 1 can reach all other vertices. The code used to generate these instances is available in the repository.
  • Road Graphs: 12 graphs representing road networks in different regions of the United States, obtained from the 9th DIMACS Implementation Challenge (dimacs). Edge weights correspond to the average time required to travel along the respective street. Since these are road networks, all graphs are sparse. The number of vertices and edges for each instance is reported in Table 2.
For all instances, shortest paths were computed using vertex 1 as the source. Each algorithm was executed five times independently on each instance to reduce the influence of external factors such as operating system variations or timing fluctuations. The average execution time, measured in milliseconds, was used as the main performance metric. The results are discussed in Section 6.
6 Results
6.1 Sparse Random Graphs
Table 1 reports execution times for sparse random graphs with mean out-degree of 3. Experiments were conducted on graphs with up to $2^{25}$ vertices, the maximum size that could be accommodated within the available memory. Attempts to process a graph with $2^{26}$ vertices resulted in memory exhaustion. The scaling of average execution time (on a logarithmic scale) across graph sizes is illustrated in Figure 1.
Figure 1: Execution time for random sparse graphs of increasing size, on a log-log scale.
Figure 2 illustrates the ratio of execution times between the BMSSP and Dijkstra algorithms on random graphs. For the smallest graph size ($128$ vertices), BMSSP is substantially slower, measuring approximately $24$ times slower than Dijkstra. However, as the graph size increases, this performance disparity significantly narrows, with the ratio dropping to approximately $3$ for some graphs. This decreasing ratio indicates that BMSSP becomes relatively more efficient as the input size grows. Overall, across all random graph instances, BMSSP maintains an average execution time approximately four times higher than Dijkstra.
Table 1: Execution times (in milliseconds) of our implementations of Dijkstra and Duan et al. algorithms on random sparse graphs.
Instance $n$ $m$ Time (ms) Ratio
dljkstra1959note duan2025breaking
random128D3 $2^{7}$ $3\times 2^{7}$ $0.006\pm 0.003$ $0.148\pm 0.012$ 24.667
random256D3 $2^{8}$ $3\times 2^{8}$ $0.013\pm 0.004$ $0.163\pm 0.010$ 12.538
random512D3 $2^{9}$ $3\times 2^{9}$ $036\pm 0.005 0.0$$087\pm 0.011$ 7.194
random1024D3 $2^{10}$ $3\times 2^{10}$ $0.087\pm 0.006$ $0.575\pm 0.021$ 6.609
random2048D3 $2^{11}$ $3\times 2^{11}$ $0.192\pm 0.006$ $1.175\pm 0.028$ 6.120
random4096D3 $2^{12}$ $3\times 2^{12}$ $0.431\pm 0.017$ $2.244\pm 0.038$ 5.206
random8192D3 $2^{13}$ $3\times 2^{13}$ $1.036\pm 0.081$ $634\pm 0.058 4.4$73
random16384D3 $2^{14}$ $3\times 2^{14}$ $2.199\pm 0.040$ $8.639\pm 0.410$ 3.929
random32768D3 $2^{15}$ $3\times 2^{15}$ $4.769\pm 0.068$ $20.455\pm 0.539$ 4.289
random65536D3 $2^{16}$ $3\times 2^{16}$ $11.237\pm 0.143$ $44.527\pm 0.413$ 3.963
random131072D3 $2^{17}$ $3\times 2^{17}$ $28.035\pm 0.654$ $95.711\pm 1.126$ 3.414
random262144D3 $2^{18}$ $3\times 2^{18}$ $73.609\pm 0.903$ $237.379\pm 2.058$ 3.225
random524288D3 $2^{19}$ $3\times 2^{19}$ $174.674\pm 2.412$ $631.082\pm 1.475$ 3.613
random1048576D3 $2^{20}$ $3\times 2^{20}$ $418.830\pm 2.058$ $1450.863\pm 0.565$ 3.464
random2097152D3 $2^{21}$ $3\times 2^{21}$ $946.968\pm 7.659$ $3226.209\pm 8.977$ 3.407
random4194304D3 $2^{22}$ $3\times 2^{22}$ $2012.560\pm 3.322$ $8218.205\pm 13.655$ 4.083
random8388608D3 $2^{23}$ $3\times 2^{23}$ $4494.410\pm 5.503$ $15646.055\pm 11.004$ 3.481
random16777216D3 $2^{24}$ $3\times 2^{24}$ $9998.167\pm 6.375$ $34104.239\pm 188.505$ 3.411
random33554432D3 $2^{25}$ $3\times 2^{25}$ $21765.062\pm 14.522$ $85382.069\pm 49.101$ 3.923
To check if the implementation matches the expected behavior of the algorithm, we can use the asymptotic complexities of the algorithm to simulate the same ratio as the graph size increases:
$f(n)=\frac{\text{Time(BMSSP)}}{\text{Time(Dijkstra)}}=\frac{n\log^{2/3}n}{n\log n}=\log^{-1/3}n.$
Figure 3 illustrates the expected ratio based on asymptotic analysis, and it is consistent with the ratio obtained experimentally.
Figure 2: Experimental ratio of BMSSP to Dijkstra execution times on random sparse graphs.
Figure 3: Theoretical ratio of BMSSP to Dijkstra execution times based on asymptotic analysis.
6.2 Road Graphs
Table 2 reports execution times for the 12 road network instances, and Figure 4 illustrates the growth of execution time with increasing graph size. The results show that Dijkstra’s algorithm consistently outperforms the BMSSP algorithm by an average factor of four across all twelve road network instances. The performance ratio remains remarkably stable, ranging only from $3.67$ to $4.89$, despite graph sizes varying over two orders of magnitude.
The observed four-fold performance ratio is similar to the ratio previously measured on random graphs. Dijkstra maintains a clear advantage; the performance gap seems stable as the graph size increases.
Table 2: Execution times (in milliseconds) of our implementations of Dijkstra and Duan et al. algorithms on 12 USA road network instances.
Instance $n$ (approx.) $m$ (approx.) Time (ms) Ratio
dljkstra1959note duan2025breaking
New York City $2.6\times 10^{5}$ $7.3\times 10^{5}$ $33.694\pm 0.266$ $138.444\pm 0.731$ 4.109
San Francisco Bay Area $3.2\times 10^{5}$ $8.0\times 10^{5}$ $39.672\pm 0.535$ $163.748\pm 1.228$ 4.128
Colorado $4.3\times 10^{5}$ $1.0\times 10^{6}$ $51.731\pm 0.356$ $253.033\pm 1.556$ 4.891
Florida $1.0\times 10^{6}$ $2.7\times 10^{6}$ $126.690\pm 1.733$ $546.35\pm 5.189$ 4.313
Northwest USA $1.2\times 10^{6}$ $2.8\times 10^{6}$ $153.513\pm 1.658$ $621.822\pm 2.873$ 4.051
Northeast USA $1.5\times 10^{6}$ $3.8\times 10^{6}$ $228.126\pm 1.654$ $836.506\pm 3.755$ 3.667
California and Nevada $1.8\times 10^{6}$ $4.6\times 10^{6}$ $258.069\pm 2.587$ $1006.176\pm 2.887$ 3.899
Great Lakes $2.7\times 10^{6}$ $6.8\times 10^{6}$ $378.239\pm 3.004$ $1626.529\pm 3.147$ 4.300
Eastern USA $3.5\times 10^{6}$ $8.7\times 10^{6}$ $541.000\pm 3.923$ $2413.533\pm 4.204$ 4.461
Western USA $6.2\times 10^{6}$ $1.5\times 10^{7}$ $1012.382\pm 2.023$ $549.724\pm 9.840 4.4$94
Central USA $1.4\times 10^{7}$ $4\times 10^{7} 33$$355.727\pm 12.259$ $14104.518\pm 21.586$ 4.203
Full USA $2.3\times 10^{7}$ $5.8\times 10^{7}$ $4566.797\pm 4.672$ $21261.355\pm 10.304$ 4.656
Figure 4: Execution time versus graph size for the USA road network instances, plotted on log-log scale.
6.3 General analysis
After running the tests, it was observed that the classical algorithm of dljkstra1959note achieved superior performance in all evaluated instances. Although, in theory, the algorithm proposed by duan2025breaking has a better asymptotic upper bound, its implementation is considerably more complex, involving multiple recursive calls and nuances that introduce a large constant factor in the running time. In practice, this overhead causes the algorithm to perform worse than Dijkstra’s algorithm, which is simple, direct, and highly efficient—especially when using, for example, a binary heap structure implemented directly in an array.
If we assume that the constant of our BMSSP implementation is four times higher than Dijkstra’s ($c_{2}/c_{1}=4$), a ratio derived from the road network experiment, then the same analysis used in Section 3 predicts that our implementation will only be faster when the graph size surpasses $10^{19}$.
These results indicate that the theoretical improvement obtained through asymptotic analysis does not necessarily translate into practical gains, due to the large hidden constants and structural complexity of the new algorithm. In most real-world environments, using Dijkstra’s algorithm is therefore substantially more efficient.
7 Concluding Remarks
We presented an experimental analysis comparing Dijkstra’s classical algorithm (dljkstra1959note) with the recent algorithm by duan2025breaking, which establishes the best known deterministic asymptotic SSSP upper bound in the comparison-addition model.
Our experiments show that Dijkstra’s algorithm remains the preferred choice for graphs with up to ten million vertices, covering most practical applications. Furthermore, we predict that Dijkstra’s algorithm should be the preferred choice for graphs significantly larger than that.
Dijkstra’s algorithm remains the standard for real-world graph problems. However, the work of duan2025breaking illustrates that improvements may be achievable, inspiring the exploration of algorithms that combine strong asymptotic performance with low practical overhead.
Several avenues for further investigation remain open. First, we plan to refine the BMSSP implementation by replacing its current linear-time selection routine (blum_time_1973) with a more efficient deterministic selection algorithm that offers better constant factors (alexandrescu2017fast). This modification may reduce the algorithm’s overhead and narrow the performance gap observed in our experiments.
Additionally, we will expand our experimental analysis in several directions. We intend to evaluate both algorithms on structured grid graphs, on which the Bellman-Ford algorithm has previously been shown to outperform Dijkstra (cherkassky1996shortest). We also plan to conduct systematic comparisons between expected-case and worst-case performance, implement a Fibonacci heap version of Dijkstra’s algorithm as an additional baseline, and test the impact of the degree-normalization preprocessing step required by the duan2025breaking algorithm to ensure constant out-degree.
Note
This is a work in progress.